A certain electric motor draws a current of

74 A

when connected to 120 V.

(a) What is the motor's power consumption?
W

(b) How much energy does it use during

3 h

of operation? Express the answer in joules and in kilowatt-hours.

J
kWh
Answer #1

a] Motor's power consumption = Voltage*current = 120*74 = 8880 W

b] Energy in J = power*time =  8880 W * 3h =  8880 W * 3*3600s

= 95.9*10^6 J

Energy = power*time = 8880 W * 3h

= 8.88 kW*3h

= 26.64 kWh